# A note on Gurzadyan's theorem

### 26 Mar 2026 v1

@@ ABSTRACT
The issue and proof of Gurzadyan's theorem are presented in a concise manner which avoids tedious and unnecessary calculations that would only mask what is essential. Its goal is to provide a good mathematical and physical understanding on the theorem, making you want to know more about the subject.
@@ 1 Newton's shell theorem
Consider masses uniformly distributed on a sphere $S$ at an areal density of $m$. The sphere is centered on point $C$ and has radius $b$. From superposition principle, the overall gravitational potential and gravitational field at a point $P$ are, respectively,
@@
$u(P,Q)$ being a function of the positions of point $P$ and of point $Q$, the latter being on the sphere, and given by $u(P,Q)=G/(PQ)$ from Newton's law of Gravity ($G$ is the gravitational constant). Using Gauss's theorem and taking into account the spherical symmetry of the distribution, it is easy to find the following result, first published by Newton in 1687 [1], and refered as Newton's shell theorem :
$\bullet$ The distribution acts on a test mass at a point $P$ outside the sphere as if all its mass $M=4\pi b^{2}m$ were concentrated at its center $C$, i.e. ${\bm{G}}(P,C)=-GM\displaystyle{{\bm{CP}}\over{CP^{3}}}$  ;
@@
As done in Ref. [2], it is relevant to ask if this result is a direct consequence of Newton's gravitational law (inverse-square law for the field). Without presupposing the form of $u(P,Q)$, let's then write the potential as
@@
Defining $u_{P}(Q)=u(P,Q)$, it becomes apparent that $\bar{u}_{P}(C)$ is, mathematically, the mean value of the function $u_{P}(Q)$ at point $C$, when this function is integrated over the sphere $S(C,b)$. Under what conditions would we have $\bar{u}_{P}(C)=u_{P}(C)$  ? This happens in two particular cases. One is trivial : when $u(P,Q)$ is constant. The other appears when $PC\gg b$, because then the distribution is seen from point $P$ as a point mass $M$ located at $C$. However, this does not provide any information regarding the law of forces. Is there a solution for any finite $b$  ? The answer is given by a well-known mathematical theorem (the ``Mean value theorem", see e.g. Ref. [3]) stating that the equality $\bar{u}_{P}(C)=u_{P}(C)$ is an exclusive property of harmonic functions :
@@
$\Delta_{C}$ being the Laplacian operator with respect to the coordinates of point $C$. Of course, taking the physicist's point of view, we assume that all mathematical operations necessary for a physical description are allowed. In appendix B is given a presentation of this theorem for an euclidean space of dimension $n\geq 3$, using the appropriate Green function and Dirac distribution.

### 27 Mar 2026 v2

@@ ABSTRACT
The issue and proof of Gurzadyan's theorem are presented concisely, avoiding tedious and unnecessary calculations that would mask what is essential. The goal is to provide a good mathematical and physical understanding of the theorem, making you want to learn more about its use in cosmology.
@@ 1 Newton's shell theorem
We consider masses uniformly distributed on a sphere $S$ at an areal density of $m$. The sphere is centered on point $C$ and has radius $b$. From the superposition principle, the overall gravitational potential and gravitational field at point $P$ are, respectively,
@@
$u(P,Q)$ being a function of both positions of point $P$ and point $Q$, the latter being on the sphere, and given by $u(P,Q)=G/(PQ)$ from Newton's law of gravity ($G$ is the gravitational constant). Using Gauss's theorem and considering the spherical symmetry of the distribution, it is easy to find the following result, first published by Newton in 1687 [1], and referred to as Newton's shell theorem :
$\bullet$ The distribution acts on a test mass at a point $P$ outside the sphere as if its mass $M=4\pi b^{2}m$ was concentrated at its center $C$, that is, ${\bm{G}}(P,C)=-GM\displaystyle{{\bm{CP}}\over{CP^{3}}}$  ;
@@
As done in Ref. [2], it is relevant to ask whether this result is a direct consequence of Newton's gravitational law (the inverse square law for the field). Without presupposing the form of $u(P,Q)$, we write the potential as :
@@
Defining $u_{P}(Q)=u(P,Q)$, it becomes apparent that $\bar{u}_{P}(C)$ is the mean value of the function $u_{P}(Q)$ at point $C$, when this function is integrated over the sphere $S(C,b)$. Under what conditions would we have $\bar{u}_{P}(C)=u_{P}(C)$  ? This happens in two particular cases. One is trivial : when $u(P,Q)$ is constant. The other appears when $PC\gg b$, because the distribution is then seen from point $P$ as a point mass $M$ located at $C$. However, this does not provide any information regarding the law of forces. Is there a solution for any finite $b$  ? The answer is given by a well-known mathematical theorem (the ``Mean value theorem", see e.g. Ref. [3]) stating that the equality $\bar{u}_{P}(C)=u_{P}(C)$ is an exclusive property of harmonic functions :
@@
$\Delta_{C}$ being the Laplacian operator with respect to the coordinates of point $C$. From the physicist's perspective, we assume that all mathematical operations necessary for a physical description are allowed. In appendix B is given a presentation of this theorem for an euclidean space of dimension $n\geq 3$, using the appropriate Green function and Dirac distribution.